OpenStudy (anonymous):

Solve sin(9x) + sin(7x) for all values of x

OpenStudy (agent0smith):

You can't solve something if it's not an equation...

OpenStudy (anonymous):

what to solve sin(9x)+sin(7x)=0?

OpenStudy (anonymous):

\(\sin(9x) + \sin(7x)\) is just simply \(\sin(9x) + \sin(7x)\)

OpenStudy (anonymous):

It should be sin(9x) + sin(7x) = 0

OpenStudy (anonymous):

\[\sin(p+q)=\sin(p)\cos(q)+\cos(p)\sin(q)\]\[\sin(p-q)=\sin(p)\cos(q)-\cos(p)\sin(q)\]\[\sin(p+q)+\sin(p-q)=2\sin(p)\cos(q)\]\[p+q=9x\]\[p-q=7x\] \[p=8x\]\[q=x\]Then the equation:\[\sin(9x)+\sin(7x)=2\sin(8x)\cos(x)=0\] You only need to find the x´s that make \[\sin(8x)=0\rightarrow x=k\frac{ \pi }{ 8 }\] or\[\cos(x)=0\rightarrow x=(2k-1)\frac{ \pi }{ 2 }\] where k=0,1,2.....

OpenStudy (anonymous):

The idea is that whenever you have an equation built with sin and cos, the best thing is to convert it into a product

OpenStudy (agent0smith):

Good answer, but does it find all solutions? http://www.wolframalpha.com/input/?i=sin%289x%29+%2B+sin%287x%29+%3D+0

OpenStudy (agent0smith):

Compared to http://www.wolframalpha.com/input/?i=2sin%288x%29cos%28x%29%3D0

OpenStudy (agent0smith):

Seems odd that wolfram even shows an equivalent form of it as sin(9x)+sin(7x) in the second link...

OpenStudy (anonymous):

agent0smith You are right, I should have said: k=0, +/-1, +/-2,... (must be my positive thinking...:)

OpenStudy (agent0smith):

Still, it's odd that wolfram displays the solutions so differently for two equivalent functions... and look at the number line at the bottom of each link, the solution sets are not the same (the original form has far more solutions than the modified form).

OpenStudy (anonymous):

well the solutions given by Wolfram for cos(x)=0 is \[x=\pi n-\frac{ \pi }{ 2 }=(2n-1)\frac{ \pi }{ 2 }\]which is the same as mine. I will analyze the first link because it seems quite an obscure method

OpenStudy (agent0smith):

Yeah I know your solution is the same as the second link. It's the first that seems odd.

OpenStudy (anonymous):

unnecessarily complicated I would say (which proves there is a machine behind...)

OpenStudy (agent0smith):

I think maybe it just doesn't show them all in the second :/

OpenStudy (anonymous):

First time I see Wolfram. Fishy thing, solutions are supposed to be periodic and this is not what they look like in second link